OpenStudy (anonymous):

Find the equation, in slope- intercept form, of the line passing through the specified points with the specified slopes.

OpenStudy (anonymous):

\[Point (0,-3) slope of \left(\begin{matrix}3 \\ 4\end{matrix}\right)\]

OpenStudy (anonymous):

y - - 3 = 3/4 ( x - 0) y + 3 = 3/4x y = 3/4x - 3

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

welcome!